• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name brads

    Log entry time 18:26:33 on April15,2008

    Entry number 232243

    This entry is a followup to: 231610

    Followups:

    keyword=Chamber2 U13 follow-up

    There was still a problem -- the U13 signals look the same as they did with the original cosmics run (see referenced halog entry).

    I traced the ribbon cable from the A/D card at U13 over to the level translator panel and confirmed it was good. I double checked the labels on the cable that runs from the level translator to the TDC and they appear correct (but I did not test the physical cable).

    Here is the cable run:

    Chamber2 (U13) -> Level translator IN (cable labelled 13-4T) : Level translator OUT (cable 2U13L) -> TDC slot 20, ch 64--79 (cable labelled 2U13F)

    I elected to replace the A/D card at U13 with a brand new one. The previous card was labelled "Damaged?" and put in the cabinet. Run 3107 (cosmics) was started with the new card in place.


    FIGURE 1